OpenStudy (anonymous):

Geometry help! the ratio of students who love geometry to students who just like geometry is 47 to 3. If there are 188 students who love geometry how many just like geometry?

OpenStudy (anonymous):

um i think u would do 188/47=4 so 3x4=12

OpenStudy (anonymous):

47:3 is the ratio. 47+3=50 total. You can convert it to a percent. 3/50=0.06=6%. Now, just find 6% of 188.